OpenStudy (anonymous):

find v w where v = 3i + j and w = i +3j

OpenStudy (anonymous):

v 'cross' w ? v 'dot' w?

OpenStudy (amoodarya):

ARE THEY COMPLEX NUMBER? OR VECTORS?

OpenStudy (anonymous):

i and j... i bet vectors...

OpenStudy (amoodarya):

I agree with you DemolisionWolf

OpenStudy (isaiah.feynman):

@scooper04 please we'd like you to reply

OpenStudy (anonymous):

sorry v dot x

OpenStudy (anonymous):

v 'dot' w means you multiply the coefficients of the i componenets together and multiply the coefficients of the j components together. v = 3i + j w = i +3j v 'dot' w = (3)(1)i + (1)(3)j

OpenStudy (isaiah.feynman):

The dot product of two vectors a and b is given by \[a_{x}b_{x}+a_{y}b_{y} +...+a_{n}b_{n}\]

OpenStudy (anonymous):

so the answer 3i + 3j

OpenStudy (isaiah.feynman):

Yes!

OpenStudy (anonymous):

ok what do i do with this one. find u(vdotw) where u = 2i - j, v=3i+j and w=i+4j?

OpenStudy (anonymous):

and you add the coefficients now. v 'dot' w = (3)(1)i + (1)(3)j = 3+3 = 6 a dot product will return a single value. where as a cross product will return a vector

OpenStudy (anonymous):

u=2i - j, v=3i+j and w=i+4j u(vdotw) so do the dot product for vectors V and W first..

OpenStudy (anonymous):

so the answer for the first one is 6? not 3i +3j?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

when you put this: u(vdotw) do you mean: u * (v 'dot' w) or u 'dot' (v 'dot' w)?

OpenStudy (anonymous):

u(v dot/times w)

OpenStudy (anonymous):

here is a page: http://mathinsight.org/dot_product_examples

OpenStudy (anonymous):

i'm not sure what to do with the u on this new question... I think the dot product of V and W will become a scalar that we just multiply the u vector by

OpenStudy (anonymous):

ya, I was right in what I was thinking for the second question, just checked

OpenStudy (anonymous):

so what does it look like im confused. im not good with this stuff and thursday is my last class.

OpenStudy (anonymous):

first do the dot product for vector V and W because they are inside the ( )

OpenStudy (anonymous):

that will give you a sinlgle number. multiply that number to each term in the u vector , and ur dun

OpenStudy (anonymous):

so (3)(4)i * (4)(3)j ?

OpenStudy (anonymous):

u=2i - j, v=3i+j and w=i+4j u(vdotw) = 2i - j (3*1 + 1*4)

OpenStudy (anonymous):

I gotta take off, but do you see where the "3*1 + 1*4" comes from? the coefficients of the vectors V and W, take the i's times the i's coefficients, then the j's times the j's coefficients.

OpenStudy (anonymous):

yeah i got ya

OpenStudy (anonymous):

so it the answer 1

OpenStudy (anonymous):

or 2 or neither

OpenStudy (anonymous):

cool, so then u(vdotw) = 14i - 7j

OpenStudy (anonymous):

sorry, open study crashed on me